# Joint Beamforming and Position Optimization for Fluid RIS-aided ISAC Systems

### v2

@@ III-D Optimization of Transmit Beamformer $\mathbf{W}$
$\mu^{(t)}$ $=\mu^{(t-1)}+\gamma\left(\overline{\mathbf{w}}^{\mathrm{T}(t)}\overline{\mathbf{w}}^{(t)}-P_{t}\right).$ (33b)
@@
Specifically, one can solve the subproblem (33a) by quasi-Newton method [42], which is implemented by the fminunc tool. This process is terminated when $\left|\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right|<\epsilon_{0}$ is satisfied. It is noted that the penalty factor $\gamma$ gradually increases among the iterations for fast convergence. After obtaining $\overline{\mathbf{w}}^{\star}$, it is converted to the complex matrix form and reshaped as $\mathbf{W}^{\star}$.
@@ III-E Optimization of fRIS Element Positions $\mathbf{p}$
In this subsection, we focus on the optimization of the fRIS element position $\mathbf{p}$ with other variables being fixed. First, the problem (19) can be simplified by omitting the unrelated terms and constraints:
@@
$\mathrm{s.t.}$ $~{}\mathbf{p}_{n}\in\mathcal{A},\forall n=1,\cdots,N,$ (34b)
$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}\geq\Delta D,\forall n,n^{\prime}=1,\cdots,N,n\neq n^{\prime}.$ (34c)
As mentioned in section II-A3, the channel $\mathbf{G}$ and $\mathbf{H}_{rc}$ are determined by the fRIS element positions. It is challenging to solve the problem (34) owing to two main reasons. First, the objective function is a complicated composite function concerning the variable $\mathbf{p}$. Since the fRIS element positions affect both the incident channel $\mathbf{G}$ and the reflective channel $\mathbf{H}_{rc}$, there appear quadratic and quartic forms simultaneously in the outer function. In the interior function, the variable $\mathbf{p}$ is contained in the complex exponential function of an array manifold, shown in (8). On the other hand, the constraint (34c) takes the form of “Norm $\geq$ Constant”, rendering it non-convex.
To explicitly observe $\mathbf{p}$ in (34a), we process the outer function by substituting $\mathbf{G}$ in (10) and $\mathbf{H}_{rc}$ in (11) to (34a) and integrating the unrelated symbols of $\mathbf{p}$, which is given by the following proposition.
Proposition 1. After omitting the constants, the objective function (34a) can be rewritten as
@@
Proof: Detailed derivations are provided in Appendix A.
However, the variable in (III-E) is contained in the array manifold, while there still exist $K$ quartic function terms, a quadratic function term and a linear function item with respect to the array manifold. Fortunately, it can be proved in the following proposition that the quartic, quadratic and linear items in (III-E) can be reduced to the first-order items and the difference of the first-order items if we only focus on optimizing the position of one element at a time.
Proposition 2. When considering the position of $n$-th element $\mathbf{p}_{n}$ while fixing others, the objective function $\widetilde{f}_{0}$ in (III-E) can be further simplified as
@@
$~{}~{}+\sum_{k=1}^{K}\sum_{i=1,i\neq n}^{N}\widetilde{\nu}_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right)$
@@
$~{}~{}-\sum_{k=1}^{K}\overline{\nu}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right).$ (36)
@@
In the first term, we denote $\xi_{n}$ and $\rho_{n}$ as the phase and amplitude of $n$-th entry of ${}\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\nu_{n}\!=\!2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is the constant. The symbol $d_{r,n}(\mathbf{p}_{n})$ is $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})\!+\!p_{y,n}\sin(\psi_{r})$. In the second term, $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\!=\!d_{r,n}(\mathbf{p}_{n})\!-\!d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\xi}_{i,n,k}$ is the phase of a constant $\widetilde{c}_{i,n,k}$. Meanwhile, we denote $\widetilde{\nu}_{k}=2(1\!-\!\alpha)c_{0}\zeta_{k}$. In the last term, we represent $\overline{\xi}_{n,k}$ as the phase of $s_{n,k}$ and $\overline{\nu}_{n,k}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\overline{\rho}_{n,k}$.
Proof: Detailed derivations and symbol explanations are shown in Appendix B.
@@
Observing the objective function $\widetilde{f}_{1}$, it contains cosine functions, which are neither convex nor concave with respect to $\mathbf{p}_{n}$. To address this problem, we employ the MM framework, where surrogate functions that locally upper bound the objective function $\widetilde{f}_{1}$ are constructed. Specifically, the constructions of the surrogate function is given by the following lemma.
@@
Lemma 1. [43] Considering a function $\widetilde{f}$ with $\mathbf{x}$ being the variable, its second-order Taylor expansion at the given point $\mathbf{x}_{0}$ can be given as
$\widetilde{f}(\mathbf{x})\approx\widetilde{f}(\mathbf{x}_{0})$ $+(\nabla\widetilde{f}(\mathbf{x}_{0}))^{\mathrm{T}}(\mathbf{x}-\mathbf{x}_{0})$
$+\frac{1}{2}(\mathbf{x}-\mathbf{x}_{0})^{\mathrm{T}}\nabla^{2}\widetilde{f}(\mathbf{x}_{0})(\mathbf{x}-\mathbf{x}_{0}),$ (37)
@@
where $\nabla\widetilde{f}(\mathbf{x})$ is the gradient vector of $\widetilde{f}(\mathbf{x})$, while $\nabla^{2}\widetilde{f}(\mathbf{x}_{0})$ represents the Hessian matrix of the function.
@@
By selecting a parameter $\delta$ satisfying $\delta\mathbf{I}\succeq\nabla^{2}\widetilde{f}(\mathbf{x}_{0})$, the function $\widetilde{f}(\mathbf{x})$ can be locally upper-bounded by
@@
$\!\widetilde{f}(\mathbf{x})\leq\widetilde{f}(\mathbf{x}_{0})\!+\!(\nabla\widetilde{f}(\mathbf{x}_{0}))^{\mathrm{T}}(\mathbf{x}-\mathbf{x}_{0})\!+\!\frac{\delta}{2}\left\|(\mathbf{x}-\mathbf{x}_{0})\right\|^{2}.$ (38)
@@
In particular, the parameter $\delta$ can be chosen as the Frobenius norm of the Hessian matrix $\nabla^{2}f(\mathbf{x}_{0})$.
@@
According to the Lemma 1, we separate $\mathbf{p}_{n}$ from the cosine functions and utilize the second-order Taylor expansion as the surrogate function. To construct the surrogate function, the gradient vector and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ should be determined. Specifically, the gradient and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ are given by the following theorem.
@@
Theorem 1. According to the linear property of derivative operation, the gradient of $\widetilde{f}_{1}(\mathbf{p}_{n})$ is the sum of the derivatives of its three components, which can be expressed as
@@
For the Hessian matrix of $\widetilde{f}_{1}$, it can also be given as
@@
According to Lemma 1 and Theorem 1, the equivalent objective function ultimately be given as
@@
where the constants are neglected and $\delta_{n}=\|\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})\|^{2}_{\mathrm{F}}$. The $\mathbf{p}_{n}^{(t-1)}$ represents the output position of $n$-th element from the last round optimization. Apparently, the function (41) is a convex quadratic function with respect to $\mathbf{p}_{n}$.
Apart from the objective function, the constraint (34c) remains non-convex in the current form. To deal with it, an effective approach is to relax the constraint (34c) with the first-order Taylor expansion. Specifically, the first-order Taylor expansion of the left term in (34c) can be given as
$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}$ $\geq\|\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}$
$~{}~{}~{}+\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}$ (42a)
@@
which is affine regarding $\mathbf{p}_{n}$. Therefore, the problem for optimizing the position of the $n$-th element is given as
$\min_{\mathbf{p}_{n}}$ $~{}(\nabla\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)}))^{\mathrm{T}}(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\!\frac{\delta_{n}}{2}\left\|(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\right\|^{2}$ (44a)
$\mathrm{s.t.}$ $~{}\mathbf{p}_{n}\in\mathcal{A},$ (44b)
$\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}\!\geq\!\Delta D,\forall n\!\neq\!n^{\prime}.$ (44c)
The problem (44) is a typical convex quadratic problem, which can be solved by using quadprog [44, 45]. Besides, we adopt circle packing scheme [46] for the initialization of fRIS element positions, where the initial positions are sufficiently separated. The details of the proposed algorithm are summarized in Algorithm 1.
@@
0:    Channel parameters, transmit signal, fRIS phase shift matrix, system parameters, desired reflective signal and communication symbols;
1:  Initialize the element positions $\mathbf{p}^{(0)}$ with the circle packing scheme;

### v3

@@ III-D Optimization of Transmit Waveform $\mathbf{x}$
$\mathbf{b}_{2}$ $=\alpha\left(\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{s}_{r}\right)+(1\!-\!\alpha)\omega\left(\mathbf{H}_{c}^{\mathrm{H}}\mathbf{s}_{c}\right).$ (26b)
Notably, when $\boldsymbol{\Theta}$ is an ideal modulus-one diagonal matrix, $\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ in (26b) can be reduced to $\mathbf{G}^{\mathrm{H}}\mathbf{G}$ due to $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$. After this transformation, it is seen that (25) is a convex quadratic function. For (23b), it can be re-expressed as the inner product of $\mathbf{x}$, that is, $\mathbf{x}^{\mathrm{H}}\mathbf{x}\leq P_{t}$. In this context, we consider full power usage, namely, $\mathbf{x}^{\mathrm{H}}\mathbf{x}=P_{t}$. Therefore, (23) is rewritten as
$\min_{\mathbf{x}}$ $~\mathbf{x}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{x}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{x}\}~~~~\mathrm{s.t.}~\mathbf{x}^{\mathrm{H}}\mathbf{x}=P_{t},$ (27)
which can be addressed by ALM [13, 3]. Specifically, we first employ the complex-real conversion by separating the real part and the imaginary part. Thus, (27) is converted to its real form as
$\min_{\overline{\mathbf{x}}}$ $~\mathring{f}(\overline{\mathbf{x}})=\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{A}}_{2}\overline{\mathbf{x}}-2\overline{\mathbf{b}}_{2}^{\mathrm{T}}\overline{\mathbf{x}}~~~~\mathrm{s.t.}~\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{x}}=P_{t},$ (28)
where $\overline{\mathbf{x}}=\left[\mathfrak{R}\{\mathbf{x^{\mathrm{T}}}\},\mathfrak{I}\{\mathbf{x^{\mathrm{T}}}\}\right]^{\mathrm{T}}$, $\overline{\mathbf{b}}_{2}=\left[\mathfrak{R}\{\mathbf{x^{\mathrm{T}}}\},\mathfrak{I}\{\mathbf{x^{\mathrm{T}}}\}\right]^{\mathrm{T}}$ and $\overline{\mathbf{A}}_{2}=\left[\mathfrak{R}\{\mathbf{A}_{2}\},-\mathfrak{I}\{\mathbf{A}_{2}\};\mathfrak{I}\{\mathbf{A}_{2}\},\mathfrak{R}\{\mathbf{A}_{2}\}\right]$. Accordingly, the augmented Lagrangian function of (28) can be written as
$\mathcal{L}\triangleq\mathring{f}(\overline{\mathbf{x}})+\mu\left(\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{x}}-P_{t}\right)+\frac{\gamma}{2}\left(\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{x}}-P_{t}\right)^{2},$ (29)
in which $\mu$ is the Lagrange multiplier and $\gamma$ denotes the penalty factor. Then, we tackle (29) using the following iteration procedure:
$\overline{\mathbf{x}}^{(t)}$ $=\arg\min_{\overline{\mathbf{x}}}\mathcal{L}(\overline{\mathbf{x}},\mu^{(t-1)}),$ (30a)
$\mu^{(t)}$ $=\mu^{(t-1)}+\gamma\left(\overline{\mathbf{x}}^{\mathrm{T}(t)}\overline{\mathbf{x}}^{(t)}-P_{t}\right).$ (30b)
@@
Specifically, we can solve (30a) via the quasi-Newton method [3], which is implemented by the fminunc tool. This process is terminated when $\left|\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{x}}-P_{t}\right|<\epsilon_{0}$ is satisfied.
@@ III-E Optimization of fRIS Element Positions $\mathbf{p}$
In this subsection, we focus on the optimization of the fRIS element position $\mathbf{p}$. First, (18) is simplified by omitting the unrelated terms and constraints:
@@
[…]
Inspecting (31), $\mathbf{G}$ and $\mathbf{H}_{rc}$ are determined by the variable $\mathbf{p}$, and (31a) is a complicated composite function, where quadratic and quartic forms exist simultaneously in the outer function. In the interior function, $\mathbf{p}$ is contained in the complex exponential function, shown in (7).
To explicitly observe $\mathbf{p}$ in (31a), we substitute $\mathbf{G}$ in (9) and $\mathbf{H}_{rc}$ in (10) to (31a) and integrate the unrelated symbols of $\mathbf{p}$, which is given by the following proposition.
Proposition 1. After omitting the constants, (31a) is rewritten as
@@
Proof: The proof is provided in Appendix A.
However, $\mathbf{p}$ in (III-E) is contained in the array manifold, while there exist $K$ quartic terms, a quadratic term, and a linear item w.r.t. the array manifold. Fortunately, it is shown in the following proposition that these items can be reduced to the first-order items and the difference of the first-order items if we focus on optimizing the position of one element at a time.
Proposition 2. When considering the position of the $n$-th element $\mathbf{p}_{n}$ while fixing others, $\widetilde{f}_{0}$ in (III-E) is simplified as
@@
$~~+\sum_{k=1}^{K}\sum_{i=1,i\neq n}^{N}\widetilde{\nu}\zeta_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right)$
@@
$~~-\sum_{k=1}^{K}\overline{\nu}~\overline{\rho}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right).$ (33)
@@
In the first term, we denote $\xi_{n}$ and $\rho_{n}$ as the phase and amplitude of the $n$-th entry of $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\nu_{n}\!=\!2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is the constant. The symbol of $d_{r,n}(\mathbf{p}_{n})$ is defined as $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})\!+\!p_{y,n}\sin(\psi_{r})$. In the second term, $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\!=\!d_{r,n}(\mathbf{p}_{n})\!-\!d_{c,k,n}(\mathbf{p}_{n})$, while $\widetilde{\xi}_{i,n,k}$ is the phase of a constant $\widetilde{c}_{i,n,k}$. Meanwhile, we denote $\widetilde{\nu}=2(1\!-\!\alpha)c_{0}$. In the last term, we represent $\overline{\xi}_{n,k}$ as the phase of $s_{n,k}$, while $\overline{\nu}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}$.
Proof: The proof is provided in Appendix B.
Remark 3: The derivations in Propositions 1 and 2 rely on structural assumptions of both the fRIS and the channel model. Specifically, the fRIS is required to be passive and single-connected, i.e., $|\boldsymbol{\Theta}_{i,i}|=1$. Furthermore, the channels are considered as LoS channels, which is reasonable in practical scenarios. The RIS is strategically deployed to establish virtual LoS links, for example, to bypass obstacles or extend coverage.
@@
Observing $\widetilde{f}_{1}$ in (III-E), it contains cosine functions, which are neither convex nor concave w.r.t. $\mathbf{p}_{n}$. To address this problem, we employ the MM framework, where surrogate functions that upper bound $\widetilde{f}_{1}$ are constructed. Specifically, the construction of the surrogate functions is given by the following lemma.
@@
Lemma 1. [28] Considering a function $\widetilde{f}$ with $\mathbf{z}$ being the variable, its second-order Taylor expansion at $\mathbf{z}_{0}$ is given by
$\widetilde{f}(\mathbf{z})\approx\widetilde{f}(\mathbf{z}_{0})$ $+(\nabla\widetilde{f}(\mathbf{z}_{0}))^{\mathrm{T}}(\mathbf{z}-\mathbf{z}_{0})$
$+\frac{1}{2}(\mathbf{z}-\mathbf{z}_{0})^{\mathrm{T}}\nabla^{2}\widetilde{f}(\mathbf{z}_{0})(\mathbf{z}-\mathbf{z}_{0}),$ (34)
@@
where $\nabla\widetilde{f}(\mathbf{z})$ is the gradient vector of $\widetilde{f}(\mathbf{z})$, while $\nabla^{2}\widetilde{f}(\mathbf{z})$ represents the Hessian matrix of $\widetilde{f}(\mathbf{z})$.
@@
By selecting a parameter $\delta$ satisfying $\delta\mathbf{I}\succeq\nabla^{2}\widetilde{f}(\mathbf{z}_{0})$, the function $\widetilde{f}(\mathbf{z})$ can be upper-bounded by
@@
$\!\widetilde{f}(\mathbf{z})\leq\widetilde{f}(\mathbf{z}_{0})\!+\!(\nabla\widetilde{f}(\mathbf{z}_{0}))^{\mathrm{T}}(\mathbf{z}-\mathbf{z}_{0})\!+\!\frac{\delta}{2}\left\|(\mathbf{z}-\mathbf{z}_{0})\right\|^{2}.$ (35)
@@
In particular, $\delta$ is chosen as the Frobenius norm of $\nabla^{2}\widetilde{f}(\mathbf{z}_{0})$.
@@
According to Lemma 1, we separate $\mathbf{p}_{n}$ from the cosine functions and utilize the second-order Taylor expansion as the surrogate function. To construct the surrogate function, the gradient vector and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ should be determined, which are given by the following theorem.
@@
Theorem 1. According to the linear property of derivative operation, the gradient of $\widetilde{f}_{1}(\mathbf{p}_{n})$ is the sum of the derivatives of its three components, which is expressed as
@@
For the Hessian matrix of $\widetilde{f}_{1}$, it is given by
@@
According to Lemma 1 and Theorem 1, the relaxed objective function is given by
@@
where the constants are neglected, while $\delta_{n}=\|\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})\|^{2}_{\mathrm{F}}$. The $\mathbf{p}_{n}^{(t-1)}$ represents the position of the $n$-th element from the previous optimization result. Apparently, (38) is a convex quadratic function w.r.t. $\mathbf{p}_{n}$.
Remark 4: The construction of (38) guarantees that the surrogate and the original function match in both value and gradient at the current point $\mathbf{p}_{n}^{(t-1)}$, ensuring local approximation. In particular, the approximation error can be characterized as
error $\approx\frac{1}{2}\Delta\mathbf{p}_{n}^{\mathrm{T}}\left(\delta_{n}\mathbf{I}-\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)})\right)\Delta\mathbf{p}_{n},$ (39)
where $\Delta\mathbf{p}_{n}=\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)}$. The upper-bounding strategy ensures that the surrogate becomes increasingly tight during convergence. Nevertheless, since the original problem is non-convex, the MM algorithm is only guaranteed to converge to a stationary point, which is generally a local optimum.
Apart from the objective function, (14f) remains non-convex. To deal with it, we relax (14f) with the first-order Taylor expansion. Specifically, the first-order Taylor expansion of the left term in (14f) is given by
$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}$ $\geq\|\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}+\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}$ (40a)
@@
which is affine regarding $\mathbf{p}_{n}$. Notably, this method relaxes the non-convex feasible region to an inner half-space feasible region. Since part of the feasible region is discarded, the optimization may only find a suboptimal solution. Nonetheless, such approximation ensures algorithmic stability and is commonly accepted in MM-based optimizations. Therefore, the problem for optimizing the position of the $n$-th element is given by
$\!\!\!\!\min_{\mathbf{p}_{n}}~\nabla\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)})^{\mathrm{T}}(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})+\!\frac{\delta_{n}}{2}\left\|(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})\right\|^{2}$ (42a)
[…]
which is a typical convex quadratic problem, and solved by using quadprog [37]. Besides, we adopt the circle packing scheme [38] for the initialization of fRIS element positions, where the initial positions are sufficiently separated. This scheme guarantees minimum spacing and provides sufficient mobility space for subsequent iterations to avoid collision. The details of the proposed algorithm are summarized in Algorithm 1.
Remark 5: Element repositioning may introduce additional overhead as compared to the traditional RIS. On the one hand, element displacement consumes energy, which mainly stems from the step motor drive or the port switching. On the other hand, element mobility adds computational complexity to position optimizations. Nonetheless, the extra spatial DoFs of fRIS allow the use of fewer elements, keeping the total computational burden reasonable.
@@
0:   Channel parameters, transmit signal, fRIS phase shift matrix, reference signal, and communication symbols;
1: Initialize $\mathbf{p}^{(0)}$ with the circle packing scheme;